1/2

Caroline Keppel,  Caroline Keppel,
Treu und herzinniglich  Robin Adair

tradukita de Wilhelm Gerhard

Robin Adair

Treu und herzinniglich, Robin Adair, tausendmal grüss´ ich dich, Robin Adair!

What’s this dull town to me Robin’s not near What was’t I wish’d to see What wish’d to hear

Hab´ ich doch manche Nacht schlummerlos hingebracht, immer an dich gedacht, Robin Adair!

Where all the joy and mirth Made this town heaven on earth Oh, they’re all fled with thee Robin Adair

Dort an dem Klippenhang, Robin Adair, rief ich oft still und bang: Robin Adair!

What made th’ assembly shine Robin Adair What made the ball sae fine Robin was there

Fort von dem wilden Meer falsch ist es, liebeleer, macht nur das Herze schwer, Robin Adair!

What when the play was o’er What made my heart so sore Oh, it was parting with Robin Adair

Mancher wohl warb um mich, Robin Adair! Treu aber liebt’ ich dich, Robin Adair!

But now thou’rt cold to me Robin Adair But now thou’rt cold to me Robin Adair

Mögen sie andre frei’n will ja nur dir allein Leben und Liebe weih’n, Robin Adair!

Yet he I loved so well

Still in my heart shall dwell Oh, I can ne’er forget Robin Adair

…  …

2/2

Traduko de la Angla poemo “Robin Adair” de Caroline Keppel (∗1734 – †1769) en la Germanan de Wilhelm Gerhard (∗1780-11-29 – †1858-10-

02) en 1826.

Arg-1025-2079 (2014-05-27 15:42:39)

Tiu ĉi kanto-teksto troviĝas en http: // www. volksliederarchiv. de/ text2030. html . Pri la tradukinto vidu la retejon http: // de. wikipedia. org/ wiki/ Wilhelm_Gerhard .

Verkinto de tiu ĉi Angla poemo estas Caroline Keppel (∗1734 – †1769).

Arg-1025-2077 (2014-05-27 15:55:44)

La teksto de tiu ĉi kanto troviĝas en http: // www. traditionalmusic. co. uk/ folk-song-lyrics/ Robin_Adair. htm . Pri Robert Adair vidu la retejon http: // en. wikipedia. org/ wiki/ Robert_Adair_( politician) . Legu ankaŭ: http: // www. lookandlearn. com/ blog/ 24 82/ lady-caroline-keppels-sad-love-song-immortalised-young-robin-adair/ . Vidu ankaŭ: http: // www. bartleby. com/ 360/ 3/ 74. html . La kanto estas aŭdebla en http: // www. youtube. com/ watch? v= 51cI0_exnLU .

http://www.poezio.net/version?poem-id=1025&version-id=2079,2077